OpenStudy (anonymous):

Let f(x) = 9x – 2 and g(x) = –x + 3. Find f(g(x)). –9x + 27 –9x + 25 –9x + 5 –9x – 2

OpenStudy (anonymous):

f(x) = 9x - 2 g(x) = -x + 3 f(g(x)) = 9(-x + 3) - 2 = -9x + 25